### v1

@@ 1 Introduction
In medical image analysis, numerous machine learning models and explainable AI (xAI) methods have been proposed [wang2022trends, shi2023mapping, frasca2024explainable, bhati2024survey, muhammad2024unveiling]. While model performance is typically well evaluated, other aspects, such as the correctness of model reasoning, that is, whether a model makes the right decision for the right reasons, are often insufficiently assessed [rudin2019stop]. Even for xAI methods that are explicitly designed to improve interpretability, systematic evaluation remains a major challenge [muhammad2024unveiling]. A key reason for this limitation is the lack of comprehensive ground truth information. In addition to annotations for the target classes, the evaluation of xAI methods also requires ground truth for the visual explanations at the sample level. Such annotations are rare, particularly in the medical domain where dataset sizes are inherently limited.
@@ 2.1.1 Image Generation
Each image in the FunnyNodules dataset is synthetically generated through a parameterized algorithm that constructs an abstract nodule as a grayscale image, see samples in Table 2. The function models nodules as elliptical shapes whose geometry, boundary, and intensity are determined by the six attributes. Spiculation is simulated via angular contour perturbations, edge sharpness by Gaussian blurring, and the optional central structure by adding a small textured subregion. Random perturbations of rotation and background noise ensure slight natural variation across samples while preserving exact attribute control.
@@ 2.2 Customizability
The dataset framework is highly customizable, allowing systematic investigation of the impact of various factors on model performance. Target definitions can follow simple linear rules or be based on difficult attribute correlated conditions, enabling analysis of model behavior under different levels of task complexity. The set of visual attributes can be varied in number, type, and scale. Experiments can be conducted with or without explicitly defined regions of interest (ROIs), and the presence or absence of background introduces additional complexity in nodule detection and segmentation. Furthermore, image size can be adjusted to approximate realistic scaling, rather than relying on interpolation. Finally, input channels can be selected according to the simulated application domain, e.g., grayscale for chest X-ray or CT images and RGB for dermatological images, providing flexibility for diverse imaging tasks. While these examples illustrate key customization options, the framework is flexible and can accommodate additional variations.
@@ 3 Evaluation Methods
We used various models for our experiments, including standard models and prototype-based approaches. We trained ResNet-50 [He_2016_CVPR] and DenseNet-121 [Huang_2017_CVPR] in a multitask classification setting, where attributes and target predictions are concurrently predicted in the final layer. Models with hierarchical structures, such as Proto-Caps, HierViT, and Concept Bottleneck Networks (joint setting), naturally reflect the attribute-to-target relationship. We note that the models were not extensively optimized for the dataset, because the goal of this section is to demonstrate how FunnyNodules can be used for evaluation rather than to compare model performance.
@@ 3.1 Evaluation of a Model’s Reasoning
fig:attribute_oneAttributeChangeEffectonTarget shows the mean target values for each attribute alongside the model predictions, which enables an assessment of how consistently the model follows the ground truth trends.
@@ 3.4 Assessing Prototype Reasoning Correctness
Table 4: Prototype Correctness is an important measure of the truthfulness of a models’ prototype reasoning.
@@ 4 Discussion and Conclusion
In this work, we introduce the FunnyNodules dataset, which provides extensive opportunities for evaluating diverse aspects of model behavior and explainability methods, with a particular focus on attribute-based reasoning models. The fully controlled parametric synthesis of attribute characteristics enables the creation of multiple types of ground truth annotations, such as attribute-level scores and region-of-interest masks. Moreover, the framework’s customizability allows full control over the relationship between attributes and target classes, making it possible to simulate different levels of reasoning complexity and decision rules. We also provided example analyses highlighting the dataset’s versatility for evaluating and interpreting AI models.

### v2

@@ 1 Introduction
In medical image analysis, numerous machine learning models and explainable AI (xAI) methods have been proposed [wang2022trends, shi2023mapping, frasca2024explainable, bhati2024survey, muhammad2024unveiling]. While model performance is typically well evaluated, other aspects, such as the correctness of model reasoning, that is, whether a model makes the right decision for the right reasons, are often insufficiently assessed [rudin2019stop]. Even for xAI methods that are explicitly designed to improve interpretability, systematic evaluation remains a major challenge [muhammad2024unveiling]. A key reason for this limitation is the lack of comprehensive ground truth information. In addition to annotations for the target classes, the evaluation of xAI methods also requires ground truth for the visual explanations at the sample level. Such annotations are rare, particularly in the medical domain where dataset sizes are inherently limited [LITJENS201760].
@@ 2.1.1 Image Generation
Each image in the FunnyNodules dataset is synthetically generated through a parameterized algorithm that constructs an abstract nodule as a grayscale image, see samples in Table 2. The function models nodules as elliptical shapes whose geometry, boundary, and intensity are determined by the six attributes. Spiculation is simulated via angular contour perturbations, edge sharpness by Gaussian blurring, and the optional internal structure by adding a small textured subregion. Random perturbations of rotation and background noise ensure slight natural variation across samples while preserving exact attribute control.
@@ 2.2 Customizability
The dataset framework is highly customizable, allowing systematic investigation of the impact of various factors on model performance. Target definitions can follow simple linear rules or be based on difficult attribute correlated conditions, enabling analysis of model behavior under different levels of task complexity. The set of visual attributes can be varied in number, type, and scale. Experiments can be conducted with explicitly defined regions of interest (ROIs), and the presence of background introduces additional complexity in nodule detection and segmentation. Furthermore, image size can be adjusted to approximate realistic scaling, rather than relying on interpolation. Finally, input channels can be selected according to the simulated application domain, e.g., grayscale for chest X-ray or CT images and RGB for dermatological images, providing flexibility for diverse imaging tasks. While these examples illustrate key customization options, the framework is flexible and can accommodate additional variations.
@@ 3 Evaluation Methods
We used various models for our experiments, including standard models and prototype-based approaches. We trained ResNet-50 [He_2016_CVPR] and DenseNet-121 [Huang_2017_CVPR] in a multitask classification setting, where attributes and target predictions are concurrently predicted in the final layer. Models with hierarchical structures, such as Proto-Caps [Gallee:peerj:proto], HierViT [Gallee:arXiv:2025:Hierarchical], and Concept Bottleneck Model (joint setting) [pmlr-v119-koh20a], naturally reflect the attribute-to-target relationship. We note that the models were not extensively optimized for the dataset, because the goal of this section is to demonstrate how FunnyNodules can be used for evaluation rather than to compare model performance.
@@ 3.1 Evaluation of a Model’s Reasoning
fig:attribute_oneAttributeChangeEffectonTarget shows the mean target values for each attribute alongside the model predictions, which enables an assessment of how consistently the models follow the ground truth trends.
@@ 3.4 Assessing Prototype Reasoning Correctness
Table 4: Prototype Correctness is an important measure of the truthfulness of a models’ prototype reasoning. [train=1800, val=200, test=500]
@@ 4 Discussion and Conclusion
In this work, we introduce the FunnyNodules dataset, which provides extensive opportunities for evaluating diverse aspects of model behavior and explainability methods, with a particular focus on attribute-based reasoning models. The fully controlled parametric synthesis of attribute characteristics enables the creation of multiple types of ground truth annotations, such as attribute-level scores and region-of-interest masks. Moreover, the framework’s high customizability allows precise specification of how attributes relate to target classes, making it possible to simulate different levels of reasoning complexity and decision rules. We also provided example analyses highlighting the dataset’s versatility for evaluating and interpreting AI models.